# The Winner of the NFL Draft is Not Necessarily Cursed

### v1

@@ 2.3 Surplus value
Massey and Thaler, (2013), an economics paper at heart, are interested in surplus value, or the difference between first contract performance value and cost. Surplus value accounts for the varying cost of each draft pick (recall the cost curve from Figure 1). Formally, denoting the first contract compensation of draft pick $x$ by ${\mathrm{cost}}(x)$ and its performance outcome by $Y$, surplus value is $S=Y-{\mathrm{cost}}(x)$. Massey and Thaler, (2013) model expected surplus value ${\mathbb{E}}[S|x]$. In the previous sections, we considered nonlinear transformations of performance value that upweighted eliteness. Here, we extend this approach to surplus value, exploring nonlinear transformations of surplus, $v(x)\propto{\mathbb{E}}[g(S)|x]$.
In Figure 7 we visualize expected surplus value (black dashed line) and the probability that surplus value exceeds a threshold (colored lines). Recall the loser’s curse from Massey and Thaler, (2013) and Section 1: according to expected surplus value, last year’s worst teams who pick at the top of the draft have less valuable picks than better teams who pick later in the first round. We visualize this by the large spike in the black dashed line. If you value a pick by the probability it results in a large surplus, the effect size of the loser’s curse diminishes and the surplus peak shifts earlier in the draft.
Figure 7: Surplus value relative to the first pick ($y$-axis) as a function of draft position $x$. The colored lines are right tail probability curves proportional to ${\mathbb{P}}(S>r|x)$ and the dashed black line is proportional to expected surplus value ${\mathbb{E}}[S|x]$.
@@ 2.4 Adjusting for position
The performance $Y$ of a player drafted at position $x$ depends on his position, particularly whether he is a quarterback. Quarterbacks command the largest contracts in the NFL and have by far the highest wins above replacement according to PFF WAR (Eager,, 2020). Formally, quarterbacks have a much fatter right tail than other positions (see Figure 19 in Appendix D.2). This is because they have a much higher mean and variance than other positions (see Figure 18 in Appendix D.2). Beyond quarterback, draft curves should also potentially differ across other positions. Previous research suggests it is more difficult to forecast the value of certain positions than other positions. Also, certain positions are inherently more valuable than others (Cole,, 2023; Baldwin,, 2024). Consequently, it is important to adjust for position (${\mathsf{pos}}$), or at least for quarterbacks (${\mathsf{qb}}$), in modeling the density of performance outcomes.
Accordingly, in Appendix D.2 we estimate the conditional density ${\mathbb{P}}(Y|x,{\mathsf{pos}})$ using a Bayesian hierarchical Beta regression model. It is crucial to shrink positional estimates towards a common mean since there are so few datapoints for each position. The standard errors of the positional conditional densities (and conditional means) are huge and need to be accounted for. We don’t find large or significant differences across non-quarterback positions (see Figure 18). But, quarterbacks have significantly higher conditional mean and variance curves (and hence fatter right tails) than non-quarterbacks. Thus, in formulating draft trade value curves, we consider just quarterbacks and non-quarterbacks. We define the non-quarterback conditional density by averaging over all the non-quarterback positions, ${\mathbb{P}}(Y|x,{\mathsf{not\ qb}}):=\frac{1}{\#{\mathsf{pos}}-1}\sum_{k\neq{\mathsf{qb}}}{\mathbb{P}}(Y|x,{\mathsf{pos}}=k)$.
In Figure 8 we visualize performance value curves for quarterbacks and non-quarterbacks. We see the same general trend from Figure 4: higher eliteness cutoffs correpond to steeper curves. Further, the higher the cutoff of eliteness, the less a non-${\mathsf{qb}}$ pick is worth relative to a first pick ${\mathsf{qb}}$. For instance, the probability that the second contract of a non-${\mathsf{qb}}$ drafted first exceeds $15\%$ of the cap is about $10\%$ of that of a ${\mathsf{qb}}$ (see the purple lines). In contrast, the expected performance value of a non-${\mathsf{qb}}$ drafted first is about $60\%$ of that of a ${\mathsf{qb}}$ (see the dashed black lines). This makes sense: a non-${\mathsf{qb}}$ drafted first will on average provide less value than a ${\mathsf{qb}}$ drafted first, but nowhere near as much less value than if we value performance by eliteness. For large cutoffs of eliteness, the value of later non-${\mathsf{qb}}$ picks quickly converges to $0\%$ of a first pick ${\mathsf{qb}}$. This makes sense: just about the only way to get an extremely elite, far-right tail performance outcome is from a ${\mathsf{qb}}$.
@@
Figure 8: Performance value relative to the first ${\mathsf{qb}}$ pick ($y$-axis) as a function of draft position $x$ and position (facet). The colored lines are right tail probability curves proportional to ${\mathbb{P}}(Y>r|x,{\mathsf{pos}})$ and the dashed black line is proportional to expected performance value ${\mathbb{E}}[Y|x,{\mathsf{pos}}]$.
In Figure 9 we visualize surplus value curves for quarterbacks and non-quarterbacks. According to expected surplus value, there is a loser’s curse for ${\mathsf{qb}}$s and non-${\mathsf{qb}}$s, as seen by the hump in the dashed black line: last year’s worst teams who pick at the top of the draft have less valuable picks than better teams who pick later in the first round. If you are interested in an elite ${\mathsf{qb}}$, however, there is no loser’s curse: the surplus peak all but disappears for right tail probability surplus curves. The far-right tails of performance $Y$ are so fat for top pick ${\mathsf{qb}}$’s that even the high cost of top picks doesn’t massively diminish their value.333 See Figure 20 in Appendix D.2 to visualize the right tails of the conditional surplus densities for ${\mathsf{qb}}$s and ${\mathsf{not\ qb}}$s. For non-quarterbacks, the loser’s curse remains, even if you are interested in eliteness. Surplus value peaks in the middle of the first round and decays across the draft. This makes sense: early picks’ compensation are priced for the ultra high mean and variance of quarterbacks. On average, this price becomes too high for other positions.
@@
Figure 9: Surplus value relative to the first ${\mathsf{qb}}$ pick ($y$-axis) as a function of draft position $x$ and position (facet). The colored lines are right tail probability curves proportional to ${\mathbb{P}}(S>r|x,{\mathsf{pos}})$ and the dashed black line is proportional to expected surplus value ${\mathbb{E}}[S|x,{\mathsf{pos}}]$.
@@ 3 Discussion
Traditional NFL draft value curves are expected value curves: the value of a draft position is proportional to expected performance or surplus value. Massey and Thaler, (2013), who first constructed these curves, had two main findings. First, they coined the loser’s curse: top draft picks, belonging to last year’s worst teams, produce less surplus value on average than draft picks later in the first round that belong to better teams. Second, the draft value curve implied by the trade market is much steeper than expected performance and surplus value curves. Thus, according to those expected value curves, general managers trade suboptimally: they don’t trade down often enough, and they often overpay when they trade up.
Massey and Thaler’s analysis from over a decade ago completely replicates to recent data (2013-2023). Even though their work has been public and widely known for a long time, general managers’ draft behavior has remained largely unchanged. Why? It may be that general managers are stubborn and not wise. Alternatively, there is a sound economic reason why average future player value (at a draft position) doesn’t determine trade value. We are partial to the latter explanation.
The goal is to win a Super Bowl, or to be the best team of 32 teams. To win a Super Bowl, we suspect that elite players, and particularly elite quarterbacks, are exponentially more valuable than average players. Thus, we construct draft value curves that value eliteness. The most intuitive example is right tail probability: the value of a draft position is proportional to the probability that the performance or surplus of a drafted player exceeds a threshold. We find that a steep right tail probability curve nearly matches the curve implied by the trade market, suggesting that general managers may want elite players, not average players. They may be willing to pay quite a lot for draft picks that are likely to be elite NFL performers. Then, we adjust for position, creating separate expected value and right tail probability draft curves for quarterbacks and non-quarterbacks. We find that, if you are looking for an elite quarterback, there is no loser’s curse. If you are looking for elite talent, picks from lower draft picks aren’t worth much and the high cost of a top draft pick is justified.
The right objective function is Super Bowl win probability: the value of a draft position should be proportional to the probability that a drafted player increases the probability you win the Super Bowl in the next $t$ years. We do not solve this problem because it is extremely difficult. How the addition of a draft pick changes a team’s Super Bowl win probability depends on the current composition of his team (in particular, projected current and future talent level at each position), injury risk, salary cap considerations, and projected outcomes of players in the draft. This results in an extremely fascinating and complicated problem, which we leave to future work.
The curves detailed in this work are simple proxies of Super Bowl win probability. Which proxy is right to use depends on the aforementioned factors (team composition, salary cap, and projected outcomes of players in the draft) and a team’s goals. In general, teams should trade draft picks based on some combination of eliteness and expected value. For example, consider the general manager of a below average team without an elite quarterback (think the 2023 Bears) whose goal is to win the Super Bowl sometime soon. Intuitively, drafting a median or even a good player barely increases the probability of achieving his goal, but drafting a potentially highly elite player (such as Caleb Williams) significantly increases it. Thus, we recommend using a steep trade value chart. This may explain why the Bears did not relinquish its number one pick in the 2024 draft. Now, consider the general manager of a consecutive Super Bowl winning team featuring an elite quarterback, tight end, and defensive tackle (think the 2023 Chiefs) whose goal is to win the Super Bowl sometime soon. We recommend using a shallower curve than that of the Bears. The difference in Super Bowl win probability between adding an average player and a bust is larger for the Chiefs than for the Bears. An average player can provide a real impact to the Chiefs winning another Super Bowl.
@@
Though our analysis builds upon the draft curve literature, it is not without limitations, and there are many avenues for future work. First, the methodology underlying the fitted trade market curve (the Weibull curve from Appendix C) has imperfections, which we detail in the Appendix. In particular, if the team trading down uses a different value function than the team trading up, how do we explain trading? Second, in future work, we suggest exploring curves built from the sum of performance and surplus, rather than each one individually. When you draft a player, you benefit from the value of their performance in addition to the extra cash from a cheap rookie contract. Third, a major assumption employed in this work and in previous work is additivity: the value of a bundle of draft picks $v(\{x_{j}\})$ equals its sum, $\sum_{j}v(x_{j})$. Concretely, in terms of right tail probability, we consider the value of a bundle of picks by $\sum_{j}{\mathbb{P}}(Y_{j}>r)$, though the relationship between the picks may be non-additive or something more general, like ${\mathbb{P}}(f(Y_{1},...,Y_{n})>r)$. We leave an exploration of non-additve draft curves to future work. Fourth, in this study we consider just first contract performance and surplus value. First contract value is an incomplete measure of the total value he provides to his team, as it excludes the rest of his career. Finally, we did not address a selection bias that was explored in Eager, (2024): worse teams draft players earlier in each round of the draft than better teams, and a team’s quality may influence the ultimate performance outcome of a drafted player. A more elaborate construction of NFL draft trade value curves should account for this selection bias, which we leave to future work.
@@ Acknowledgments
The authors thank Cade Massey, Christopher N. Avery, and Blake Zilberman for giving the authors a reason to explore draft value curves. We also thank Cade and Chris for sending NFL draft data for an early version of this project.
@@ Appendix C Estimating the trade market value of draft picks
Appendix D Estimating the conditional density of the outcome of a draft pick
D.1 Position-agnostic conditional density
In Figure 11 we visualize the empirical conditional density of $Y$ given $x$. The empirical densities are extremely wobbly and noisy due to limited data for each individual draft pick. But there is a general trend: we see a spike near zero that shifts upwards as draft pick increases and a right tail that shifts leftwards and gets smaller as draft pick increases. This motivates our “Bayesian spike plus Beta regression” model of ${\mathbb{P}}(Y|x)$.444 We initially tried a lone Beta regression model without the spike near zero, which less successfully fit the right tail.
Figure 11: Empirical distribution (histogram and smoothed density) of performance outcome $Y$ (second contract APY cap percentage, $x$-axis) for each draft pick $x$ (facet).
We model the spike near zero, the bust spike, using logistic regression. We let a bust be a player whose performance outcome $Y$ is smaller than some cutoff ${y_{\text{bust}}}$. We found that ${y_{\text{bust}}}=0.005$ (or $0.5\%$ of the cap) works well. We model the bust probability ${\mathrm{bp}}(x):={\mathbb{P}}(Y\leq{y_{\text{bust}}}|x)$ by
$Y\leq{y_{\text{bust}}}|x\sim{\mathrm{Bernoulli}}({\mathrm{bp}}(x)),$ (D.1)
where
${\mathrm{bp}}(x)={\mathrm{logistic}}(\alpha_{0}+\alpha_{1}\cdot x).$ (D.2)
Next, we model the right tail using Beta regression. We view the Beta distribution as a natural way to model $Y$ because it lies in $[0,1]$ and it seems reasonable when we eyeball the empirical densities (á la Figure 11). We model
$Y|x,Y>{y_{\text{bust}}}\sim{\mathrm{Beta}}(\mu(x),\phi(x)).$ (D.3)
Here, we use the mean-precision parameterization of the Beta distribution. The mean $\mu$ and precision $\phi$ are related to the traditional Beta distribution parameters $\mathrm{shape}_{1}$ and $\mathrm{shape}_{2}$ by $\mathrm{shape}_{1}=\mu\phi$ and $\mathrm{shape}_{2}=(1-\mu)\phi$. We let the mean and precision parameters vary as draft pick $x$ varies via $\mu(x)$ and $\phi(x)$ to capture that different draft picks $x$ yield different outcome distributions. The conditional mean is ${\mathbb{E}}[Y|x,Y>{y_{\text{bust}}}]=\mu(x)$ and the conditional variance is $\mathrm{Var}[Y|x,Y>{y_{\text{bust}}}]=\mu(x)(1-\mu(x))/(1+\phi(x))$. So, the precision is inversely proportional to the variance. We use spline regression to model the conditional mean,
$\mu(x)={\mathrm{logistic}}(\widetilde{x}^{\top}\beta),$ (D.4)
where $\widetilde{x}$ is a B-spline basis.555 The ${\mathsf{R}}$ code we used to generate the spline basis is splines::bs($x$, degree=3, df=4, intercept=T). We model the conditional precision by
$\phi(x)=\exp(\gamma_{0}+\gamma_{1}\cdot x).$ (D.5)
To capture uncertainty in the parameters, we use a Bayesian model, placing priors on the parameters. We put a weakly informative Normal prior with mean 0 and s.d. 10 on each of the $\alpha$, $\beta$, and $\gamma$ parameters. Finally, we assume that $Y$ is uniformly distributed within the bust spike in a way that doesn’t vary with $x$, $Y|x,Y\leq{y_{\text{bust}}}\sim{\mathrm{Unif}}[0,{y_{\text{bust}}}]$.666 We assess the reasonableness of this assumption later in this section. But this assumption is benign because when $Y\leq{y_{\text{bust}}}$, $Y$ is essentially 0 and so negligbly impacts the right tail probability calculations that form the draft value curves that we detail in this paper.
In Figure 12 we see that our estimates of the conditional mean $\mu(x)$, conditional standard deviation ${\mathrm{sd}}(x)$, and bust probability ${\mathrm{bp}}(x)$ fit the data well. In Figure 13 we see that our estimates of the right tail ${\mathbb{P}}(Y|x,Y>{y_{\text{bust}}})$ fit the data well enough. As expected, the right tail shifts leftwards as draft pick increases. In Figure 14 we visualize the conditional density ${\mathbb{P}}(Y|x)$ for values $Y$ in the right tail $Y>{y_{\text{bust}}}$. By Bayes rule, this is equal to ${\mathbb{P}}(Y|x,Y>{y_{\text{bust}}})\cdot(1-{\mathrm{bp}}(x))$. We see that the right tail, measuring the probability of drafting an elite player, shifts leftwards and decreases in size as draft pick increases. The posterior credible intervals of these density curves are fairly narrow because the information of our probability model is distilled into just a few parameters. In Figure 15 we visualize the conditional density of surplus $S=Y-{\mathrm{cost}}$. The high cost of the first few picks shifts those curves to the left relative to curves of other picks.
Figure 12: Left: our estimate of the conditional mean $x\mapsto\mu(x)={\mathbb{E}}[Y|x,Y>{y_{\text{bust}}}]$ given not a bust. Middle: our estimate of the conditional standard deviation $x\mapsto{\mathrm{sd}}(x)={\mathrm{sd}}(Y|x,Y>{y_{\text{bust}}})$ given not a bust. Right: our estimate of bust probability $x\mapsto{\mathrm{bp}}(x)={\mathbb{P}}(Y\leq{y_{\text{bust}}}|x)$. The solid black line is the posterior mean, the gray shaded region is the $95\%$ credible interval, and the black dots are the empirical values for each draft pick $x$.
Figure 13: The solid black lines are the fitted conditional density ${\mathbb{P}}(Y|x,Y>{y_{\text{bust}}})$ (the posterior mean of our model). This overlays the empirical distribution (gray bars and gray lines) for each draft pick.
Figure 14: For values $Y$ in the right tail $Y>{y_{\text{bust}}}$, the estimated conditional density ${\mathbb{P}}(Y|x)={\mathbb{P}}(Y|x,Y>{y_{\text{bust}}})\cdot(1-{\mathrm{bp}}(x))$ for various values of draft pick $x$ (color). Figure (a) displays the posterior mean density curves and (b) displays 95% credible intervals of these curves.
Figure 15: For values $Y$ in the right tail $Y>{y_{\text{bust}}}$, the estimated conditional surplus density ${\mathbb{P}}(S|x)={\mathbb{P}}(Y-{\mathrm{cost}}(x)|x,Y>{y_{\text{bust}}})\cdot(1-{\mathrm{bp}}(x))$ for various values of draft pick $x$ (color) in the first round (a) and throughout the draft (b).
Figure 16: The empirical mean (left) and standard deviation (right) of $Y$ for each draft pick $x$ conditional on being a bust $Y\leq{y_{\text{bust}}}$.
D.2 Position-specific conditional density
In this section, we model the conditional density of performance outcome $Y$ given draft pick $x$ and position ${\mathsf{pos}}$. We extend the “Bayesian spike plus Beta regression” model from Appendix D.1 to adjust for position. We use 11 position groups: quarterback (QB), wide receiver (WR), running back / fullback (RB/FB), tight end (TE), offensive tackle (OT), interior offensive lineman (IOL), interior defensive lineman (DL), edge (ED), linebacker (LB), cornerback (CB), and safety (S). We filter out all draft picks from our dataset belonging to other positions (e.g., punter, kicker, and long snapper).
It is imperative to use a Bayesian model to adjust for position, rather than fitting separate curves for each position, since there is such little data within each position group. Across each of the position groups, the median sample size in our dataset is 269 players, with a maximum of 350 and a minimum of 127. Spread across 256 total draft pick positions, this is not an abundance of data. A Bayesian hierarchical model is an excellent choice here because it allows us to capture uncertainty in position curves while also shrinking towards an overall mean.
$Y|x,{\mathsf{pos}},Y>{y_{\text{bust}}}\sim{\mathrm{Beta}}(\mu(x,{\mathsf{pos}}),\phi(x,{\mathsf{pos}})),$ (D.6)
where
$\mu(x,{\mathsf{pos}})={\mathrm{logistic}}(\widetilde{x}^{\top}\beta_{\mathsf{pos}})$ (D.7)
and
$\phi(x,{\mathsf{pos}})=\exp(\gamma_{0,{\mathsf{pos}}}+\gamma_{1,{\mathsf{pos}}}\cdot x).$ (D.8)
${\mathrm{bp}}(x,{\mathsf{pos}}):={\mathbb{P}}(Y\leq{y_{\text{bust}}}|x,{\mathsf{pos}})={\mathrm{logistic}}(\alpha_{0,{\mathsf{pos}}}+\alpha_{1,{\mathsf{pos}}}\cdot x).$ (D.9)
We shrink the position-specific coefficients $\alpha_{\ast,{\mathsf{pos}}}$, $\beta_{\ast,{\mathsf{pos}}}$, and $\gamma_{\ast,{\mathsf{pos}}}$ towards overall mean parameters $\alpha_{\ast}$, $\beta_{\ast}$, and $\gamma_{\ast}$. For example, we model the position-specific coefficients $\alpha_{0,{\mathsf{pos}}}$ as a draw from a Normally distributed prior with mean $\alpha_{0}$ and standard deviation $\tau_{\alpha_{0}}$. We put a weakly informative Normal prior with mean 0 and s.d. 10 on the prior mean parameter $\alpha_{0}$. We put a mild positive standard Normal prior on the prior standard deviation parameter $\tau_{\alpha_{0}}$. The priors and hyperpriors are similar for the other parameters.
In Figure 17a we see that our estimate of the conditional mean $\mu(x,{\mathsf{pos}})$ fits the data well. The empirical standard deviation values in Figure 17b we are much more sparse than the empirical conditional mean values in Figure 17a. This is because, for each position ${\mathsf{pos}}$ and for many draft picks $x$, just one player of position ${\mathsf{pos}}$ was drafted at $x$. Thus, there is no empirical standard deviation at many combination of $x$ and ${\mathsf{pos}}$. Nonetheless, based on the data we have observed, our estimate of the conditional standard deviation ${\mathrm{sd}}(x,{\mathsf{pos}})$ fits the data reasonably well.
Figure 17: Figure (a): our estimate of the conditional mean $\mu(x,{\mathsf{pos}})={\mathbb{E}}[Y|x,{\mathsf{pos}},Y>{y_{\text{bust}}}]$ given not a bust. Figure (b): our estimate of the conditional standard deviation ${\mathrm{sd}}(x,{\mathsf{pos}})={\mathrm{sd}}(Y|x,{\mathsf{pos}},Y>{y_{\text{bust}}})$ given not a bust. The solid black line is the posterior mean, the gray shaded region is the $95\%$ credible interval, and the black dots are the empirical values for each draft pick $x$.
In Figure 18a we see that the posterior means of the conditional mean, standard deviation, and bust probability curves vary by position. WR and OT have the highest non-quarterback conditional mean lines and RB/FB, S, LB, and TE have the lowest non-quarterback conditional mean lines. WR and ED have the highest non-quarterback conditional s.d. lines and RB/FB, S, and TE have the lowest non-quarterback conditional s.d. lines. The bust probability lines are clumped into two groups, one about $0.05$ probability higher than the other. In Figure 18b we add 95% credible intervals (the shaded regions) to these curves. The credible intervals are wide due to a dearth of data across the draft at each individual position. The non-quarterback mean and s.d. curves jumble into one big clump. Closer inspection may reveal two clumps of non-quarterback positions, one of whose curves lie slightly above the other’s, but the effect size of this difference is small. Thus, in Figure 18c we average all of the non-quarterback curves into the red curves, $\mu(x,{\mathsf{not\ qb}})=\frac{1}{\#{\mathsf{pos}}-1}\sum_{{\mathsf{pos}}\ k\neq{\mathsf{qb}}}\mu(x,{\mathsf{pos}})$ and similarly for ${\mathrm{sd}}(x,{\mathsf{not\ qb}})$ and ${\mathrm{bp}}(x,{\mathsf{not\ qb}})$. As expected, quarterbacks have a higher mean than other positions. What’s more interesting is that quarterbacks have astoundingly higher variance than other positions. This is what drives the fat far right tail of top pick quarterbacks.
Figure 18: Figure (a): The posterior mean estimates (lines) of $\mu(x,{\mathsf{pos}})$, ${\mathrm{sd}}(x,{\mathsf{pos}})$, and ${\mathrm{bp}}(x,{\mathsf{pos}})$ as a function of draft pick $x$ and position ${\mathsf{pos}}$ (color). In Figure (b) we add 95% credible intervals (the shaded regions) to these curves. Figure (c): These curves for ${\mathsf{pos}}\in\{{\mathsf{qb}},{\mathsf{not\ qb}}\}$.
In Figure 19 we visualize the conditional density ${\mathbb{P}}(Y|x,{\mathsf{pos}})$ for ${\mathsf{pos}}\in\{{\mathsf{qb}},{\mathsf{not\ qb}}\}$ (where the ${\mathsf{not\ qb}}$ density is the average of all the non-quarterback positional densities). As the draft progresses (Figure 19b), for both ${\mathsf{qb}}$’s and ${\mathsf{not\ qb}}$’s the right tail shifts leftwards and decreases in size as draft pick increases. The ${\mathsf{qb}}$ right tails are fatter than those for ${\mathsf{not\ qb}}$. For picks in the first round (Figure 19a), the ${\mathsf{qb}}$ densities have an incredibly fat right tail, significantly fatter than those of ${\mathsf{not\ qb}}$. This reflects the immense value of top pick quarterbacks: these top picks yield the highest probability of acquiring the most elite players in the game, who provide an outsize influence on winning the Super Bowl. In Figure 20 we visualize the conditional density of surplus $S=Y-{\mathrm{cost}}$ for ${\mathsf{qb}}$ and ${\mathsf{not\ qb}}$. For non-quarterbacks, the high cost of the first few picks shifts the right tails of ${\mathsf{not\ qb}}$ to the left of those of later picks, creating a non-decreasing surplus value curve even with our nonlinear performance outcome value functions (like right tail probability). For quarterbacks, even after accounting for cost the right tails of top picks remain fatter than those of later picks. This reflects the immense value of top pick ${\mathsf{qb}}$’s.
Figure 19: For values $Y$ in the right tail $Y>{y_{\text{bust}}}$, the estimated conditional density ${\mathbb{P}}(Y|x,{\mathsf{pos}})={\mathbb{P}}(Y|x,{\mathsf{pos}},Y>{y_{\text{bust}}})\cdot(1-{\mathrm{bp}}(x,{\mathsf{pos}}))$ for ${\mathsf{pos}}\in\{{\mathsf{qb}},{\mathsf{not\ qb}}\}$ (facet) and various values of draft pick $x$ (color) in the first round (a) and throughout the draft (b).
Figure 20: For values $Y$ in the right tail $Y>{y_{\text{bust}}}$, the estimated conditional surplus density ${\mathbb{P}}(S|x,{\mathsf{pos}})={\mathbb{P}}(Y-{\mathrm{cost}}(x)|x,{\mathsf{pos}},Y>{y_{\text{bust}}})\cdot(1-{\mathrm{bp}}(x,{\mathsf{pos}}))$ for ${\mathsf{pos}}\in\{{\mathsf{qb}},{\mathsf{not\ qb}}\}$ (facet) and various values of draft pick $x$ (color) in the first round (a) and throughout the draft (b).

### v2

@@ 6.2 Surplus value
Massey and Thaler, (2013) assess draft pick value through the lens of surplus value, the difference between a player’s performance during his first contract and the cost associated with his draft position (see Figure 2). Formally, denoting the first contract compensation at draft position $x$ by ${\mathrm{cost}}(x)$ and a drafted player’s performance outcome by $Y$, surplus value is defined by $S=Y-{\mathrm{cost}}(x)$. While Massey and Thaler, (2013) modeled expected surplus value, ${\mathbb{E}}[S|x]$, we consider the right tail of surplus value, ${\mathbb{P}}(S>r|x)$.
In Figure 15 we visualize expected surplus value (black dashed line) and the probability that surplus value exceeds a threshold (colored lines). The loser’s curse––top draft picks yield lower expected surplus value than those selected later in the first round––manifests as a spike in the expected surplus value curve above one during the first round. According to the probability that surplus value exceeds a high threshold, however, the effect size of the first-round spike diminishes and the surplus peak shifts earlier in the draft.
Figure 15: Value relative to the first pick ($y$-axis) as a function of draft position $x$. The long-dashed black line is the expected surplus value curve. The colored lines are proportional to right tail probability $x\mapsto{\mathbb{P}}(S>r|x)$, where $r$ is an eliteness cutoff (color). The lines are the posterior means and the shaded regions are the $95\%$ credible intervals.
@@ 6.3 Adjusting for position
The distribution of performance value varies not only by draft position but also by player position (see Section 5.2). Quarterback performance in particular exhibits a substantially fatter right tail than other positions (see Figure 12). Hence, in this section, we construct separate draft position value curves for quarterbacks and non-quarterbacks.
We begin by visualizing performance value curves in Figure 16. For ${\mathsf{pos}}\in\{{\mathsf{qb}},{\mathsf{not\ qb}}\}$, we compute expected performance value (proportional to ${\mathbb{E}}[Y|x,{\mathsf{pos}}]$) and right tail probability (proportional to ${\mathbb{P}}(Y>r|x,{\mathsf{pos}})$) from the estimated conditional density ${\mathbb{P}}(Y|x,{\mathsf{pos}})$. We compute value relative to a first pick ${\mathsf{qb}}$. The position-specific curves exhibit the same broad pattern observed in the overall draft value curves from Figure 14: right tail probability curves are steeper than expected performance value curves. As the eliteness threshold $r$ increases, right tail probability curves become progressively steeper, reinforcing the idea that teams prioritizing increasingly elite talent assign greater relative value to earlier picks.
@@
Figure 16: Value relative to the first ${\mathsf{qb}}$ pick ($y$-axis) as a function of draft position $x$ and position (facet). The dotted black line is expected performance value. The colored lines are proportional to right tail probability $x\mapsto{\mathbb{P}}(Y>r|x,{\mathsf{pos}})$, where $r$ is an eliteness cutoff (color). The lines are posterior means and the shaded regions are $95\%$ credible intervals.
In Figure 17 we visualize position-specific surplus value curves. For both quarterbacks and non-quarterbacks, expected surplus value peaks in the first round. According to the probability that surplus value exceeds a high threshold, however, there is no first-round peak for quarterbacks. This suggests that, for teams seeking elite quarterbacks, the extreme upside of top selections offsets their higher cost. In contrast, even for elite non-quarterbacks, the first-round spike persists. Surplus value for non-quarterbacks peaks in the early-to-mid first round and declines steadily throughout the draft. The right tails of top-pick non-quarterbacks are not sufficiently fat to justify the high cost of early selections, suggesting that those picks are primarily priced for the ultra-high mean and variance of quarterbacks.
@@
Figure 17: Value relative to the first ${\mathsf{qb}}$ pick ($y$-axis) as a function of draft position $x$ and position (facet). The long-dashed black line is expected surplus value. The colored lines are proportional to right tail surplus probability $x\mapsto{\mathbb{P}}(S>r|x,{\mathsf{pos}})$, where $r$ is an eliteness cutoff (color). The lines are posterior means and the shaded regions are $95\%$ credible intervals.
@@ 7 Discussion
A central premise in the judgment and decision making literature is that experts, even in high-stakes environments, exhibit systematic cognitive biases. Massey and Thaler, (2013) applied this framework to the NFL draft, concluding that, according to expected surplus value, general managers overvalue top draft picks, a finding widely cited as evidence of irrational decision making. This conclusion, however, hinges on the assumption that general managers should optimize for expected surplus value, an assumption that is neither explicitly justified nor necessarily aligned with the strategic realities of NFL team-building.
In this paper, we challenge that assumption and present an alternative framework for evaluating the value of a draft position. By estimating the full conditional density of player performance given draft position and player position, we construct draft position value curves from alternative objective functions. We focus on right tail probability––equivalently, the number of elite players expected to arise from a given pick––reflecting that general managers may be prioritizing the acquisition of elite, game-changing players. Under this perspective, top draft picks are not inefficiently overvalued but rather priced for their potential to produce transformational talent—players who can dramatically elevate a franchise’s trajectory. When valuing draft position based on right tail probability rather than expected surplus value, the first-round spike in the value of a draft position largely disappears. Teams selecting at the top of the draft retain higher relative value when their objective is to secure elite talent rather than merely optimizing expected returns. This effect is particularly pronounced for quarterbacks, whose performance distributions exhibit extremely fat right tails.
This study contributes to the judgment and decision making literature by highlighting the importance of specifying the decision-maker’s utility function before labeling behavior as irrational. If general managers value variance and right-tail outcomes rather than expected surplus, then their draft strategies, far from being biased or inefficient, may be consistent with a rational objective function tailored to the realities of NFL competition. Perhaps irrationality and biases are not as pervasive as Massey and Thaler, along with the broader judgement and decision making literature, assert. More broadly, our findings underscore the risks of applying oversimplified economic models to complex, domain-specific decision-making environments without considering alternative utility structures.
While our analysis provides a compelling alternative to the traditional expected performance and surplus value framework, it does not resolve the question of how general managers ought to value draft picks. A more comprehensive analysis would explicitly estimate the impact of draft selections on Super Bowl win probability, accounting for factors such as team composition––including projected current and future talent at each position––injury risk, salary cap constraints, and projected outcomes of available draft prospects. Prioritizing expected surplus value may be misguided, but so may prioritizing eliteness and variance; determining the optimal balance remains an open question. We suspect the best objective combines elements of both.
A broader implication of this study is the potential disconnect between formal decision models and the intuitive expertise of practitioners. NFL general managers may not be explicitly solving optimization problems or consciously weighing right-tail probabilities, yet their actions suggest an implicit alignment with valuing elite performance. This raises an important question: are expert decision-makers in other fields adept at intuitively navigating their domains in ways that are difficult for modelers to formalize? If so, then economists, statisticians, and decision scientists should approach expert behavior with greater humility before declaring it irrational. While analytical models provide valuable insights, they may overlook nuanced decision making processes that emerge from experience, tacit knowledge, and the complexity of real-world environments.
At the same time, our assumption that general managers’ primary goal is to win the Super Bowl may not be entirely correct. It is possible that NFL general managers optimize for other outcomes, such as improving their win-loss record, making the playoffs, or simply preserving their own job security. Moreover, even if their decisions align with maximizing the probability of drafting elite players, this does not necessarily imply rationality. Their choices could stem from overconfidence in their ability to identify talent, a misperception of the distribution of player outcomes, or other psychological biases. The fact that the trade market aligns with a $19.7\%$ of salary cap threshold of eliteness may be a coincidence rather than the product of deliberate strategic thinking. Future research should explore how expert intuition interacts with formal modeling across domains, identifying where models enhance decision making and where they risk oversimplifying the expertise of practitioners.
@@
Finally, we conclude with a discussion of technical limitations of our study and directions for future work. First, the methodology underlying the fitted trade market curve (the Weibull curve from Appendix C) has imperfections, which we detail in that section. In particular, if the team trading down uses a different value function than the team trading up, how do we explain trading? Second, in future work, we suggest exploring curves built from the sum of performance and surplus, rather than each one individually. When drafting a player, a general manager benefits from the value of his performance in addition to the extra cash from a cheap rookie contract. Third, a major assumption employed in this work and in previous work is additivity: the value of a bundle of draft picks $v(\{x_{j}\})$ equals its sum, $\sum_{j}v(x_{j})$. Concretely, in terms of right tail probability, we consider the value of a bundle of picks by $\sum_{j}{\mathbb{P}}(Y_{j}>r)$, though the relationship between the picks may be non-additive or something more general, like ${\mathbb{P}}(f(Y_{1},...,Y_{n})>r)$. We leave an exploration of non-additve draft curves to future work. Fourth, in this study we consider just first contract performance and surplus value. First contract value is an incomplete measure of the total value he provides to his team, as it excludes the rest of his career. Finally, we did not address a selection bias that was explored in Eager, (2024): worse teams draft players earlier in each round of the draft than better teams, and a team’s quality may influence the ultimate performance outcome of a drafted player. A more elaborate construction of NFL draft trade value curves should account for this selection bias, which we leave to future work.
@@ Acknowledgments
The authors thank Cade Massey, Christopher Avery, and Blake Zilberman for giving the authors a reason to explore draft value curves. We also thank Cade and Chris for sending NFL draft data for an early version of this project.